OpenStudy (anonymous):

By using the method of matrices, solve the following the simultaneous equation. 2x - y =4 3x - y = 5 Could someone explain in detail with the working steps?

OpenStudy (amistre64):

there are a few ways to do this with matrix manipulations. the idea tho is to establish rules used from elimination method: you can multiply rows, and add columns

OpenStudy (amistre64):

\begin{matrix} *1/2)&2&-1&|&4\\ &3&-1&|&5\\ \end{matrix} \begin{matrix} *-3)&1&-1/2&|&2\\ &3&-1&|&5\\ \end{matrix} \begin{matrix} &-3&3/2&|&-6\\ +)&3&-1&|&5\\ &-&-&-&-\\ &0&1/2&|&-1 \end{matrix} now we can reconstruct it back subbing in this new row[2] \begin{matrix} &1&-1/2&|&2\\ &0&1/2&|&-1 \end{matrix} another addition of columns and will get rid of the 1/2 parts and so on and so forth

OpenStudy (amistre64):

the easier matrix move to do is to use cramers rule: \begin{matrix} x&y&|&=\\ -&-&-&-\\ a&b&|&n\\ c&d&|&m \end{matrix} \[y=\frac{am-cn}{ad-cb}\] \[x=\frac{bm-dn}{cb-ad}\]